OpenStudy (anonymous):

Evaluate as a fraction or an integer: (3^6) (3^-3) -27 27 -1/27 1/27

OpenStudy (solomonzelman):

\[\huge\color{blue}{ a^{-b}=\frac{1}{a^b} }\]this way,\[\huge\color{blue}{ 3^{-3}=\frac{1}{3^3}=\frac{1}{27} }\]

OpenStudy (anonymous):

Fred pays 2/5 of his annual salary in federal, state, and local taxes. His annual salary is 102,000. How much of his salary does he get to keep after taxes are paid? $61,000 $72,300 $61,200 $94,900

OpenStudy (solomonzelman):

So, (for your first problem)\[\color{blue}{ 3^6 \times 3^{-3}~~~~~~is~~same~~~as,~~~~~3^6 \times\frac{1}{27} ~~~which~~is~~same~~as, }\]\[-->~~~\frac{3^6}{27}\]

OpenStudy (anonymous):

ya i get the first one now.

OpenStudy (solomonzelman):

\[\color{red}{ \frac{3^6}{3^3} ~~~->~~~3^6 \div 3^3~~~->~~~3^{(6-3)}~~~->~~~3^3~~~->~~~27. }\]

OpenStudy (anonymous):

The anwer is 27 right?

OpenStudy (anonymous):

Could you help me on the second problem?

OpenStudy (solomonzelman):

YES. Problem 2: @cliche, If he pays 2/5 of his salary he gets to keep 3/5 correct?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

We know that his ENTIRE SALARY is 102,000. if he pays 2/5, he keeps 3/5 because the entire salary is 5/5 (five fifths). So what is 3/5 of 100,200? \[\huge\color{red}{ 102,000 \times \frac{3}{5} =~~~you~~~tell~~me. }\]

OpenStudy (anonymous):

x^m *x^-n=x^(m-n)

OpenStudy (solomonzelman):

yeah, I was just showing how that works.

OpenStudy (solomonzelman):

\[\huge\color{black}{ D \times \frac{s}{v}=D \times s \div v } \]

OpenStudy (anonymous):

im confused i tried it and cam out with this 2871044762448589/46912496118441

OpenStudy (anonymous):

@SolomonZelman